OpenStudy (anonymous):

What is the slope of the line passing through the points (2,7) and (-1, 3) ?

OpenStudy (anonymous):

y2 - y1 ------ = slope x2 - x1

OpenStudy (anonymous):

So 7-3 and 2-(-1) ?

jimthompson5910 (jim_thompson5910):

keep going Luckylexi

OpenStudy (anonymous):

I got 4/-1 ? But that's not one of the answer choices.

jimthompson5910 (jim_thompson5910):

2 - (-1) is the same as 2 + 1

OpenStudy (anonymous):

So 4/3 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Can you help me with another question ?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Okaye i'll attach the file. It shows the question.

jimthompson5910 (jim_thompson5910):

ok first off, you need to find the slope of that line

OpenStudy (anonymous):

4/6 ?

jimthompson5910 (jim_thompson5910):

hmm it's a small

jimthompson5910 (jim_thompson5910):

the slope is not 4/6

OpenStudy (anonymous):

Well i'm very very confused then !

jimthompson5910 (jim_thompson5910):

remember, the slope formula is m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

Oh. Soo -4/0 ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you would only have 0 in the denominator if you had a vertical line

jimthompson5910 (jim_thompson5910):

remember, y goes on top

OpenStudy (anonymous):

4/-6 ?

jimthompson5910 (jim_thompson5910):

again, the y differences go on top of the fraction

OpenStudy (anonymous):

-6/4 ?

jimthompson5910 (jim_thompson5910):

closer, but no

jimthompson5910 (jim_thompson5910):

notice how the line has a positive slope (it goes up as you read it from left to right)

OpenStudy (anonymous):

okaye if you have (0,1) Which one is y and which one is x ?

jimthompson5910 (jim_thompson5910):

if you have the point (0,1), then x = 0 and y = 1

jimthompson5910 (jim_thompson5910):

all points are of the form (x,y)

OpenStudy (anonymous):

So -6/-4 ?

jimthompson5910 (jim_thompson5910):

now reduce that

OpenStudy (anonymous):

-3/-2 ?

jimthompson5910 (jim_thompson5910):

two negatives divide to make a _____

OpenStudy (anonymous):

positive ? so 3/2 ?

jimthompson5910 (jim_thompson5910):

perfect, the slope is 3/2

OpenStudy (anonymous):

So 3/2x+3 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

How about : The x intercept of a line is -5 and the y- intercept of a line is -2 . WHat is the equation of the line ?

OpenStudy (anonymous):

My answer is 5/2x-2 . Is that right ?

jimthompson5910 (jim_thompson5910):

that is incorrect

OpenStudy (anonymous):

Dang . What about - 5/2x-5 ?

jimthompson5910 (jim_thompson5910):

sounds like you're guessing

jimthompson5910 (jim_thompson5910):

what I recommend is you using the slope formula again

jimthompson5910 (jim_thompson5910):

the two points you will use are (-5,0) and (0,-2)

OpenStudy (anonymous):

How do i Do that if I only have two numbers ?

jimthompson5910 (jim_thompson5910):

the two intercepts mean that you have these two points on the line (-5,0) and (0,-2)

OpenStudy (anonymous):

-2/-5 ?

jimthompson5910 (jim_thompson5910):

you messed up somewhere, the overall slope should be negative

OpenStudy (anonymous):

so if they're both negative signs that means the slope will be positive ?

jimthompson5910 (jim_thompson5910):

yes, -2/-5 = 2/5 but the slope should be negative

OpenStudy (anonymous):

Well then iDk hat it could be ?

jimthompson5910 (jim_thompson5910):

it should be -2/5 so you either lost a negative sign or you introduced one in there somewhere

jimthompson5910 (jim_thompson5910):

so again, the slope is -2/5

OpenStudy (anonymous):

Okaye. I see now . I have one last one ! And i'll attach the file. Thank you soooo much !!!

jimthompson5910 (jim_thompson5910):

first off what is slope intercept form?

OpenStudy (anonymous):

-5/2 ?

jimthompson5910 (jim_thompson5910):

no in general, what is slope intercept form

OpenStudy (anonymous):

Oh y=mx+b

jimthompson5910 (jim_thompson5910):

good, so you have to turn the original equation given into that form

OpenStudy (anonymous):

So y= -5/2(x+11) ?

jimthompson5910 (jim_thompson5910):

that's not in the form y = mx + b

jimthompson5910 (jim_thompson5910):

and you simplified incorrectly

OpenStudy (anonymous):

Well how do I do that ?

jimthompson5910 (jim_thompson5910):

y - 7 = -5/2(x + 4) y - 7 = -5/2(x) - 5/2(4) y - 7 = -5/2(x) - 5/2*(4/1) y - 7 = -5/2(x) - 20/2 y - 7 = -5/2(x) - 10 I'll let you finish up

OpenStudy (anonymous):

y-7+7 = -5/2(x) -10+7 ?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

y= -5/2(x) - 3

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (anonymous):

Thank You soooo much !

jimthompson5910 (jim_thompson5910):

you're welcome